# LiRank: Industrial Large Scale Ranking Models at LinkedIn

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; August 25–29, 2024; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Information systems Recommender systemsccs: Information systems Learning to rank
@@ 2. Related Work
The use of deep neural network models in personalized recommender systems has been dominant in academia and industry since the success of the Wide&Deep model(Cheng et al., 2016) in 2016. Typically, these models consist of feature embeddings, feature selection, and feature interaction components, with much research focused on enhancing feature interactions. Tge Wide&Deep model(Cheng et al., 2016) initiated this trend by combining a generalized linear model with an MLP network. Subsequent research aimed to keep the MLP network for implicit feature interactions and replace the linear model with other modules for capturing explicit higher-order feature interactions. Examples include DeepFM(Guo et al., 2017b), which replaced the linear model with FM; deep cross network (DCN)(Wang et al., 2017) and its follow-up DCNv2(Wang et al., 2020), which introduced a cross network for high-order feature interactions; xDeepFM(Lian et al., 2018), offering compressed interaction network (CIN) for explicit vector-wise feature interactions; AutoInt(Song et al., 2018), which introduced self-attention networks for explicit feature interaction; AFN(Cheng et al., 2019), exploring adaptive-order feature interactions through a logarithmic transformation layer; and FinalMLP(Mao et al., 2023), which achieved impressive performance by combining two MLPs.
We experimented with and customized these architectures for various LinkedIn recommender tasks, with DCNv2 proving to be the most versatile. We propose enhancements to DCNv2, referred to as Residual DCN, in this paper. Additionally, we implemented a model parallelism design in TensorFlow(TF), similar to the approach proposed in the DLRM(Naumov et al., 2019a) paper, to accelerate model training with large embedding tables.
@@
While enhancing neural network predictive performance through various optimizations and architectures, the space of calibration remained relatively stable. Traditional industry-standard methods (Guo et al., 2017a) like Histogram binning, Platt Scaling, and Isotonic Regression are applied in post-processing steps after deep model training. Some research has introduced calibration-aware losses to address under/over calibration issues usually resulting in trade-off (Guo et al., 2021) or slight improved metrics (Anil et al., 2022). In §3.4 we propose an isotonic calibration layer within the deep learning model which learns to calibrate deep model scores during model training and improves model predictive accuracy significantly.
@@ 3.1. Feed Ranking Model
The primary Feed ranking model employs a point-wise ranking approach, predicting multiple action probabilities including like, comment, share, vote, and long dwell and click for each ¡member, candidate post¿ pair. These predictions are linearly combined to generate the final post score. A TF model with a multi-task learning (MTL) architecture generates these probabilities in two towers: the click tower for probabilities of click and long dwell, and contribution tower for contribution and related predictions. Both towers use the same set of dense features normalized based on their distribution(Haldar et al., 2019), and apply multiple fully-connected layers. Sparse ID embedding features (§A.1) are transformed into dense embeddings (Naumov et al., 2019b) through lookup in embedding tables of Member/Actor and Hashtag Embedding Table as in Figure  1. For reproducability in appendix in Figure 8 we provide a diagram showing how different architectures are connected together into a single model.
@@ 3.7. Member History Modeling
To model member interactions with platform content, we adopt an approach similar to (Xia et al., 2023; Chen et al., 2019). We create historical interaction sequences for each member, with item embeddings learned during optimization or via a separate model, like (Pancha et al., 2022). These item embeddings are concatenated with action embeddings and the embedding of the item currently being scored (early fusion). A two-layer Transformer-Encoder (Vaswani et al., 2017) processes this sequence, and the max-pooling token is used as a feature in the ranking model. To enhance information, we also consider the last five sequence steps, flatten and concatenate them as additional input features for the ranking model. To reduce latency, we experimented with shorter sequences and smaller feed-forward network dimensions within the Transformer. In ablation experiments in §5.2 we refer to history modeling as TransAct.
Our findings show that a two-layer transformer with a feed-forward dimension equal to half the input embedding size delivers most relevance gains. While longer sequences improve relevance metrics, the added training and serving time did not justify extended history sequences.
@@ 5.2. Feed Ranking
This methodology allows unbiased offline comparison of experimental models (Li et al., 2011). We use offline replay to assess Feed Ranking models, referred to as ’contribution’ throughout the paper (Table 6). The table illustrates the impact of various production modeling techniques on offline replay metrics, including Isotonic calibration layer, low-rank DCNv2, Residual DCN, Dense Gating, Large MLP layer, Sparse Features, MTL enhancements, TransAct, and Sparsely Gated MMoE. These techniques, listed in Table 6, are presented in chronological order of development, highlighting incremental improvements. We’ve deployed these techniques to production, and through online A/B testing, we observed a 0.5% relative increase in the number of member sessions visiting LinkedIn.
@@
Baseline -
+ 30 dimensional ID embeddings (IDs) $+1.89\%$
@@
+ Large MLP (LMLP) $+1.23\%$
+ Dense Gating (DG) +1.00%
+ Multi-task (MTL) Grouping $+0.75\%$
+ Low-rank DCNv2 (LDCNv2) $+1.26\%$
+ TransAct $+1.66\%$
+ Residual DCN (RDCN) $+2.15\%$
+ LDCNv2+LMLP+TransAct $+3.45\%$
+ RDCN+LMLP+TransAct $+3.62\%$
+ Sparsly Gated MMoE +4.14%
Table 6. Ablation study of model architecture components in Feed ranking on the relative off-policy measurement.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671561isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Neural networksccs: Information systems Recommender systemsccs: Information systems Learning to rank
@@ 2. Related Work
The use of deep neural network models in personalized recommender systems has been dominant in academia and industry since the success of the Wide&Deep model(Cheng et al., 2016) in 2016. Typically, these models consist of feature embeddings, feature selection, and feature interaction components, with much research focused on enhancing feature interactions. The Wide&Deep model(Cheng et al., 2016) initiated this trend by combining a generalized linear model with an MLP network. Subsequent research aimed to keep the MLP network for implicit feature interactions and replace the linear model with other modules for capturing explicit higher-order feature interactions. Examples include DeepFM(Guo et al., 2017b), which replaced the linear model with FM; deep cross network (DCN)(Wang et al., 2017) and its follow-up DCNv2(Wang et al., 2020), which introduced a cross network for high-order feature interactions; xDeepFM(Lian et al., 2018), offering compressed interaction network (CIN) for explicit vector-wise feature interactions; AutoInt(Song et al., 2018), which introduced self-attention networks for explicit feature interaction; AFN(Cheng et al., 2019), exploring adaptive-order feature interactions through a logarithmic transformation layer; and FinalMLP(Mao et al., 2023), which achieved impressive performance by combining two MLPs.
We experimented with and customized these architectures for various LinkedIn recommender tasks, with DCNv2 proving to be the most versatile. We propose enhancements to DCNv2, referred to as Residual DCN, in this paper. Additionally, we implemented a model parallelism design in TensorFlow(TF), similar to the approach proposed in the DLRM(Naumov et al., 2019) paper, to accelerate model training with large embedding tables.
@@
While enhancing neural network predictive performance through various optimizations and architectures, the space of calibration remained relatively stable. Traditional industry-standard methods (Guo et al., 2017a) like Histogram binning, Platt Scaling, and Isotonic Regression are applied in post-processing steps after deep model training. Some research has introduced calibration-aware losses to address under/over calibration issues usually resulting in trade-off (Guo et al., 2021; Yan et al., 2022) or slight improved metrics (Anil et al., 2022). In §3.4 we propose an isotonic calibration layer within the deep learning model which learns to calibrate deep model scores during model training and improves model predictive accuracy significantly.
@@ 3.1. Feed Ranking Model
The primary Feed ranking model employs a point-wise ranking approach, predicting multiple action probabilities including like, comment, share, vote, and long dwell and click for each ¡member, candidate post¿ pair. These predictions are linearly combined to generate the final post score. A TF model with a multi-task learning (MTL) architecture generates these probabilities in two towers: the click tower for probabilities of click and long dwell, and contribution tower for contribution and related predictions. Both towers use the same set of dense features normalized based on their distribution(Haldar et al., 2019), and apply multiple fully-connected layers. Sparse ID embedding features (§A.1) are transformed into dense embeddings (Naumov et al., 2019) through lookup in embedding tables of Member/Actor and Hashtag Embedding Table as in Figure  1. For reproducability in appendix in Figure 8 we provide a diagram showing how different architectures are connected together into a single model.
@@ 3.7. Member History Modeling
To model member interactions with platform content, we adopt an approach similar to (Xia et al., 2023; Chen et al., 2019). We create historical interaction sequences for each member, with item embeddings learned during optimization or via a separate model, like (Pancha et al., 2022). These item embeddings are concatenated with action embeddings and the embedding of the item currently being scored (early fusion). A Transformer-Encoder (Vaswani et al., 2017) processes this sequence, and the max-pooling token is used as a feature in the ranking model. To enhance information, we also consider the last five sequence steps, flatten and concatenate them as additional input features for the ranking model. From an ablation study we found that the optimal learning rate for the model with TransAct was similar to the model without TransAct. For the number of encoder layers, going from zero (just pooling) to one layer provides the largest gains, one to two layers smaller gains, and no additional gains beyond three layers. When changing the feedforward dimension as multiples of the embedding size we observe slight additional gains by going from 1/2x to 1x, 2x, and 4x. Similar trends were observed for increasing the sequence length from 25, to 50, to 100. To optimize for latency we used two encoder layers, feedforward dimension as 1/2x the embedding dimension, and sequence length 50. In ablation experiments in §5.2 we refer to history modeling as TransAct.
@@ 5.2. Feed Ranking
This methodology allows unbiased offline comparison of experimental models (Li et al., 2011). We use offline replay to assess Feed Ranking models, referred to as ’contribution’ throughout the paper (Table LABEL:table:feed_ablation_models). The table illustrates the impact of various production modeling techniques on offline replay metrics, including Isotonic calibration layer, low-rank DCNv2, Residual DCN, Dense Gating, Large MLP layer, Sparse Features, MTL enhancements, TransAct, and Sparsely Gated MMoE. These techniques, listed in Table LABEL:table:feed_ablation_models, are presented in chronological order of development, highlighting incremental improvements. We have deployed these techniques to production, and through online A/B testing, we observed a 0.5% relative increase in the number of member sessions visiting LinkedIn.
@@
Latency
CPU
Usage
Baseline
+ 30-dim ID embeddings $+1.89\%$ $+20\%$
@@
+ Large MLP $+1.23\%$ $+17\%$
+ Dense Gating +1.00%
+ Multi-task Grouping $+0.75\%$
+ Low-rank DCNv2 $+1.26\%$ $+13\%$
+ TransAct $+1.66\%$ $+52\%$ $+44\%$
+ Residual DCN $+2.15\%$ $+17\%$
+ LDCNv2+LMLP+TransAct $+3.45\%$ N/A N/A
+ RDCN+LMLP+TransAct $+3.62\%$ N/A N/A
+ Sparsly Gated MMoE +4.14% N/A N/A
Table 6. Feed ranking ablation study results. Given are percentage increases for contributions, latency, and CPU usage. A dash ”–” indicates neutrality.